OpenStudy (anonymous):

(BEST ANSWER GETS MEDAL!) Two cars leave town A at the same time and travel to town B, 480 miles away. The speed of the slower car is 20 mph less than the speed of the faster car. If the slower car arrives at town B two hours after the faster car, what is the speed of the slower car?

OpenStudy (perl):

you can use equation d = r* t car 1: d1 = r1 * t1 car 2 : d2 = r2* t2 We are given that d1 = d2 = 480 assume car 2 is slower car therefore r2 = r1 - 20 t2 = t1 + 2

OpenStudy (perl):

substituting we get 480 = r1 * t1 480 = (r1 - 20 ) ( t1 + 2)

OpenStudy (anonymous):

Ahh thank you so much! Can you help me with another problem too? :)

OpenStudy (perl):

whats the speed of the slower car?

OpenStudy (anonymous):

um not quite sure how to do that I just realized...

OpenStudy (perl):

you need to solve that system of equations you can plug the system of equations into wolfram

OpenStudy (anonymous):

okay one second